# GAPartManip: A Large-scale Part-centric Dataset for Material-Agnostic Articulated Object Manipulation

### v1

@@ V-B Actionable Pose prediction Experiments
Data Preparation. We split the dataset into training and testing sets using an approximate 7:3 ratio. Specifically, we further divide the test sets into 3 categories: seen instances, unseen but similar instances, and novel instances. We compare our methods with the following baselines:
  • GSNet (GS) [47] is a grasping pose prediction model trained on the GraspNet-1 billion [34] dataset for rigid object. We evaluate both the pre-trained model and the fine-tuned model separately.
  • Where2Act (WA)[2] is an affordance-based method for interacting with articulated objects. Unlike the original approach, we do not train a separate network for each task. As Where2act cannot generate stable grasping poses, we integrated GSNet, as referenced in [12], to enhance where2act’s capabilities to align with experimental setting.
  • EconomicGrasp (EG) [42] is also a pose prediction method for rigid objects, which includes an interactive grasp head and composite score estimation to enhance the precision of specific grasps.
Evaluation Metrics. Following [34], we utilize precision to evaluate the performance of actionable pose estimation.
$\small Precision_{\mu}=n_{suc_{\mu}}/n_{grasp}$ (3)
$Precision_{\mu}$ represents the success rate (SR) of predicted interaction poses at friction coefficient $\mu$, where $n_{grasp}$ denotes the number of predicted poses, and $n_{suc_{\mu}}$ denotes the number of successful grasps predicted under $\mu$.
Figure 6: Qualitative comparison of actionable pose prediction on synthetic data.
Results and Analysis. Our quantitative results in simulation are presented in Tab. II. Even though both GSNet and our Part-aware EcoGrasp are trained on our data, they outperform Where2Act, possibly because Where2Act struggles with cross-category and cross-action reasoning. Part-aware EcoGrasp and fine-tuned GSNet show a substantial improvement in precision compared to pretrained models. It is evident that our data significantly enhances the capability of existing methods in actionable pose estimation for articulated objects. Specifically, our dataset offers strong geometric priors for parts, enabling networks to focus more on actionable parts rather than non-actionable links. For instance, although the pre-trained EconomicGrasp in Fig. 6 generates a set of stable grasping poses, it cannot differentiate whether these poses act on actionable parts, meaning they may fail in interacting with articulated objects.
@@
* indicates that the method is trained on the GAPartManip dataset.
@@ V-C Real-World experiment
Figure 7: Qualitative Results For Real-world Manipulation. The top-15 scored actionable poses are displayed, with the red gripper representing the top-1 pose.
To validate the sim-to-real generalizability of GAPartManip, we conducte real-world experiments. We use a Franka robot arm with an Intel RealSense camera to capture depth and IR images. We compare our method with three baselines: Where2act, AO-Grasp, GSNet, and, like in V-B, We modified the Where2act interaction pipeline to finish our tasks. The experiment consists of 7 distinct instances, including StorageFurniture, Box, and Microwave, evaluating the success rate of the top-1 interactive pose for each method across open (n=14) and close (n=17) tasks. As shown in Tab. III, the overall success rate of GAPartManip is 61.29%, showcasing not only a successful transfer to the real world but also a significant performance boost compared to other methods.
Additionally, we perform ablation studies to assess how different modules affect the overall pipeline performance. As shown in Fig. 7, depth cameras yield poor depth data when faced with certain materials, significantly impacting subsequent manipulations. Our depth reconstruction module effectively addresses this issue by repairing 2D depth map, thereby enhancing the performance of subsequent modules. Similarly, as shown in Fig 7, GAPartManip tends to prioritize interactable GAParts. This part-aware capability could possibly explain why our method leads to such significant performance disparities as seen in Tab. III.
TABLE III: Real-world articulated objects manipulation results
@@
AO-Grasp 28.57 29.41 29.03
Where2act 21.42 17.64 19.35
GSNet 42.85 23.53 32.25
@@ VI Conclusions
In this paper, we build a large-scale synthetic dataset for generalizable and actionable part manipulation with material-agnostic articulated objects. Our dataset is the first large-scale, diverse in instances, categories, scenes, and materials articulated object dataset. Meanwhile, we propose an articulated object manipulation framework capable of zero-shot transfer to the real world. We conduct experiments on individual modules and real-world overall experiments, with results indicating the competitiveness of our approach. Our dataset will be released.

### v2

@@ V-B Actionable Pose Prediction Experiments
Data Preparation. We split the dataset into training and testing sets using an approximate 7:3 ratio. We further divide the testing set into 3 categories: seen instances, unseen but similar instances, and novel instances. We compare our methods with the following baselines:
  • GSNet (GS) [45] is a grasping pose prediction model trained on the GraspNet-1 billion [33] dataset for rigid objects. We evaluate both the pre-trained model and the fine-tuned model separately.
  • Where2Act (WA)[2] is an affordance-based method for interacting with articulated objects. Unlike the original approach, we do not train separate networks for each task. Since Where2Act cannot generate grasping pose prediction, we integrated GSNet, as referenced in [12], to enhance Where2Act’s capability to align with the experimental setting.
  • EconomicGrasp (EG) [40] is also a pose prediction method for rigid objects, which includes an interaction grasp head and composite score estimation to enhance the accuracy of specific grasp prediction.
Evaluation Metrics. Following [33], we utilize precision to evaluate the performance of actionable pose estimation:
$Precision_{\mu}=n_{suc_{\mu}}/n_{grasp}.$ (3)
$Precision_{\mu}$ represents the ratio of successful interaction pose prediction under the specific friction coefficient $\mu$, where $n_{grasp}$ denotes the number of predicted poses, and $n_{suc_{\mu}}$ denotes the number of successful grasps under $\mu$.
Figure 6: Qualitative Comparison of Actionable Pose Prediction in Simulation.
Results and Analysis. Our quantitative results in simulation are presented in Tab. II. Even though all fine-tuned on our dataset, both GSNet and our Part-aware EcoGrasp outperform Where2Act, possibly because Where2Act struggles with cross-category and cross-action reasoning. Our model and the fine-tuned GSNet show a substantial improvement in precision compared to the pretrained models. It is evident that our dataset significantly enhances the capability of existing methods in actionable pose estimation for articulated objects. Specifically, our dataset offers strong geometric priors for parts, enabling networks to focus more on the actionable parts rather than the non-actionable ones. For instance, although the pre-trained EconomicGrasp in Fig. 6 generates a set of stable grasping poses, it cannot differentiate whether these poses act on actionable parts, meaning they may fail in interacting with articulated objects.
@@
* indicates that the method is fine-tuned on the GAPartManip dataset.
@@ V-C Real-World experiment
Figure 7: Qualitative Results For Real-world Manipulation. The actionable poses with top scores are displayed, with the red gripper representing the top-1 pose.
To validate the sim-to-real generalizability of our novel framework, we conduct real-world experiments. We use a Franka robot arm with an Intel RealSense camera to capture depth and IR images. We compare our method with three baselines: Where2Act, AO-Grasp, GSNet, and, like in Sec. V-B, we extend the Where2Act interaction pipeline to finish our tasks. The experiments consist of 7 distinct instances, including StorageFurniture, Box, and Microwave. We evaluate the success rate of the top-1 interaction pose for each method across open (n=14) and close (n=17) tasks. As shown in Tab. III, the overall success rate of our framework is 61.29%, showcasing not only a successful transfer to the real world but also a significant performance boost compared to other methods.
Additionally, we perform ablation studies to assess how different modules affect the overall framework performance. As shown in Fig. 7, the depth camera yields poor depth maps when facing certain challenging materials, significantly impacting subsequent manipulation. Our depth reconstruction module effectively addresses this issue by recovering the depth map, thereby enhancing the performance of subsequent modules. Similarly, as illustrated in Fig. 7, Our framework, built upon our GAPartManip dataset, tends to prioritize the interaction poses on actionable parts during prediction. This part-aware capability could possibly explain why our method leads to such significant performance improvement shown in Tab. III.
TABLE III: Quantitative Results for Real-world Articulated Object Manipulation
@@
AO-Grasp [16] 28.57 29.41 29.03
Where2Act [2] 21.42 17.64 19.35
GSNet [45] 42.85 23.53 32.25
@@ VI Conclusions
In this paper, we build a large-scale, part-centric synthetic dataset for material-agnostic articulated object manipulation. Our dataset is the first dataset featuring photo-realistic material randomization and part-oriented, scene-level actionable interaction pose annotation for articulated objects. Building upon our dataset, we propose a novel articulated object manipulation framework capable of zero-shot transfer to the real world. We conduct experiments on individual modules and real-world overall experiments, with results indicating the competitiveness of our approach. Our dataset will be open-sourced.